# dsld: A Socially Relevant Tool for Teaching Statistics

### v1

@@ 4.1 Relevant Methods Provided by dsld
The dsld package provides wrappers for several functions from the faiML package (Scutari, 2023a) and for the Explicitly Deweighted Features (EDF) methods developed in Matloff and Zhang (2022). Let’s consider the fairML functions first.
FairML
Komiyama et al. (2018), Zafar et al. (2017) and Scutari (2023a) apply ridge-regression-like approaches to reducing the influence of $S$ on the predicted value, $\widehat{Y}$. All these methods are implemented in the fairML package (Scutari, 2023b), with dsld providing interfaces.
Marco Scutari’s work in the fairML package offers a statistical approach to fair machine learning, building on these previous efforts by imposing a ridge-regression-like penalty on the coefficient(s) of S. For example, let’s fit a Fair Generalized Ridge Regression model — a fair adaptation of the Generalized Linear Models, and make a prediction of recidivism in the COMPAS dataset. As an example case, take someone like observation 188 in our dataset: age 23, no prior convictions, male and Black.
cps <- compas1[,c(‘age’,‘priors_count’,‘sex’,‘two_year_recid’,‘race’)]
@@
The unfairness argument must be a value in (0,1], with values near 0 being nearly perfectly fair. This controls the Fairness-Utility Tradeoff. Here, the algorithm reports an estimated probability of recidivism of about 51%. What if we had placed no restriction on the role of race?
@@
EDF-Fair
The dsld package also includes wrappers for functions from the work on Explicitly Deweighted Features (Matloff and Zhang, 2022). This approach omits $S$ entirely and mitigates the impact of potential proxy variable through user-specified hyperparameters. The latter enables the user to select a desired point on the Fairness-Utility Tradeoff spectrum. The package offers implementations of linear/logistic regression (works like ridge regression, except with different “$\lambda$” values for each proxy variable), random forests (different proxy variables have different probabilities of being used for splitting a node), and k-nearest neighbors (k-NN) (Euclidean distance but with different weights for each proxy variable).
Let’s apply k-NN on the same example as above.
z <- dsldQeFairKNN(data=cps, yName=’two_year_recid’,
sName=‘race’, deweightPars= list(age=0.5,priors_count=0.1),
yesYVal=‘Yes’)
@@
Here, we entirely omit the race variable from our dataset and deweight the proxies ‘age’ and ‘priors count’ by 0.5 and 0.2 (a weight of 1.0 means no deweighting) to reduce their corresponding predictive power.
4.2 Assessing Fairness and Utility
One possible measure of fairness is to require that $\widehat{Y}$ and S be statistically independent (Johndrow and Lum, 2019). However, this eliminates any use of proxies and thus the ability to achieve a Fairness-Utility Tradeoff.
A broader view would be to ask that $\widehat{Y}$ and S have low correlation. See for instance Li et al. (2023), Kozodoi et al. (2022), Deho et al. (2022), Mary et al. (2019), and Baharlouei et al. (2019) apply work from Lee et al. (2022) and Roh et al. (2023).
The dsld package continues this theme:
  • Measuring Accuracy: To assess utility, we use a holdout set to compute test accuracy from predictions generated by our algorithms. In regression, accuracy is defined as mean absolute prediction error (MAPE); for classification, we use misclassification rate.
  • Measuring Fairness: The measure of fairness is the correlation between $\widehat{Y}$ and $S$, computed using the Kendall Tau statistic. This choice is preferred over the more commonly used Pearson correlation due to its flexibility with both continuous and categorical inputs. While Pearson is geared towards continuous numeric variables, Kendall Tau is also applicable for binary or ordinal integer-valued variables.
@@ 4.3 Proxy Hunting
A proxy is a variable that indirectly infers a protected attribute, potentially introducing bias in decision-making even when the protected attribute omitted from the dataset. For example, race or religion may be linked to a city or neighborhood in a city (Morse et al., 2020).
To detect potential proxies in a dataset, the dsld package provides methods, such as dsldOHunting, to help users make informed decisions. Specifically, this function calculates the Kendall Tau correlation between the sensitive feature and all other features in the dataset. Consider use of dsldOHunting on the COMPAS example presented earlier:
@@
The results suggest, say, that “age” and “prior convictions count” might be treated as proxies for being African-American, and possibly de-weighting them accordingly if one uses EDF methods.
The diverse functions offered by dsld provide a range of fair modeling techniques, allowing students to thoroughly explore the balance between fairness and accuracy in prediction tasks. Students are encouraged to apply principles of fair machine learning and illustrate the fairness-utility trade-off using different datasets.
@@ 5 Discussion
In this paper, we introduce “Data Science Looks at Discrimination” (dsld) as a powerful tool for educating students in statistics through the application of discrimination analysis and fair machine learning. The software includes a range of analytical and graphical tools that enable students to explore and visualize potential sources of bias and discrimination. The fair machine learning wrappers offer several fairness-constrained machine learning methods, facilitating the quick and seamless deployment of bias mitigation algorithms. Additionally, the accompanying Quarto notebook provides a comprehensive foundation in relevant statistical principles, featuring real-world examples that apply methods provided from dsld, and requires only a basic understanding of mathematics.
As machine learning becomes an increasingly important tool for statistical analysis and prediction, it is essential for users to understand the risks associated with its misapplication. The dsld package not only bridges the gap between traditional and fairness-constrained machine learning algorithms but also serves as an effective educational resource. It equips both current and future students with the knowledge and tools necessary to responsibly apply practical machine learning systems that balance fairness and utility.
@@
The package can be accessed here: https://github.com/matloff/dsld/tree/master. All relevant datasets, functions, and implementation details can be found through the Github repository.

### v2

@@ 4.2 Relevant Methods Provided by dsld
The dsld package provides wrappers for several functions from the fairML package (Scutari, 2023a) and for the Explicitly Deweighted Features (EDF) methods developed in Matloff and Zhang (2022). Let’s consider the fairML functions first.
Komiyama et al. (2018), Zafar et al. (2017), and Scutari (2023a) apply ridge-regression-like approaches to reduce the influence of $S$ on $\widehat{Y}$. These methods are implemented in the fairML package (Scutari, 2023b), with dsld providing interfaces.
For instance, let’s fit a Fair Generalized Ridge Regression model—a fair adaptation of generalized linear models—and predict the probability of recidivism in the COMPAS dataset. As an example, consider observation 188 in our dataset: a 23-year-old male, Black, with no prior convictions.
### subset columns
cps <- compas1[,c(‘age’, priors_count’, sex’, two_year_recid’, race’)]
newx <- cps[188,-4]
### create model
@@
The unfairness argument must be a value between (0,1], with values closer to 0 completely prioritizing fairness. This parameter controls the Fairness-Utility Tradeoff.
The algorithm reports an estimated probability of recidivism of about 51%. What if we had placed no restriction on the role of race?
@@
The dsld package also includes wrappers for functions based on EDF (Matloff and Zhang, 2022). This approach omits $S$ entirely and allows control over the degree of influence for proxy variables via user-defined ridgelike hyperparameters. Methods include linear/logistic regression (with varying “$\lambda$” values for each proxy), random forests (with different probabilities for proxy variables at each node split), and k-nearest neighbors (k-NN) (using weighted Euclidean distance to $S$ for proxies).
Let’s apply fair k-NN on the example above.
z <- dsldQeFairKNN(data=cps, yName=‘two_year_recid’, sName= race’, deweightPars=list(age=0.5,priors_count=0.1), yesYVal=‘Yes’)
@@
The probability of recidivism is 52%. In this case, race is completely excluded from our model and we control the effect of proxies using the deweightPars argument for “age” and “priors count” with weights of 0.5 and 0.1, respectively (a weight of 1.0 indicates no deweighting). This raises the question: how do we identify such proxy variables?
@@ 4.3 Proxy Hunting
A proxy is a variable that indirectly represents a protected attribute, potentially introducing bias in decision-making even when the protected attribute is not explicitly used. For example, race or religion may be linked to a city or neighborhood in a city (Morse et al., 2020).
To detect potential proxies in a dataset, the dsld package provides methods such as dsldOHunting for users to make informed decisions. This function calculates the Kendall Tau correlation between the sensitive feature and all other features in the dataset. Consider use of dsldOHunting on the COMPAS example presented earlier:
@@
Based on the correlations above, age and prior convictions might serve as proxies for being African-American, and may be deweighted accordingly when using EDF methods.
4.4 Examining the Fairness-Utility Tradeoff
The fairness-utility tradeoff indicates that prioritizing fairness often comes at the expense of predictive accuracy. Our package includes dsldUtilsFairness, which enables comparison of accuracy and fairness changes at varying levels of the unfairness and deweightPars parameter. This section provides examples using dsldQeFairKNN on the COMPAS dataset and dsldFrrm on the Law School Admissions dataset. Users can replicate this approach with other functions to efficiently conduct grid searches over different parameters and datasets.
Let’s begin with the COMPAS example. We will deweight age and prior counts with the following values: (0.01, 0.10, 0.50, 0.75, 0.99) for age, and (0.05, 0.15, 0.5, 0.95) for prior counts. dsldUtilsFairness conducts grid search across each combination of these parameters. For simplicity, the data has been pre-processed to focus on three race categories: Caucasian, African American, and Hispanic.
z <- dsldUtilsFairness(compas1, two_year_recid’, race’, dsldQeFairKNN’, deweightPars = list(age = c(0.01, 0.1, 0.5, 0.75, 0.99), priors_count = c(0.05, 0.15, 0.5, 0.95)), yesYVal = Yes’, reps = 25)
print(z)
age prior_count Accuracy black white hisp
0.01 0.05 0.2933913 0.1090585 0.09083427 0.03632643
0.10 0.05 0.2953719 0.1229806 0.10437744 0.03779181
0.50 0.05 0.2956952 0.1404892 0.11892038 0.04349329
0.75 0.05 0.2838319 0.1386847 0.11257990 0.05100799
0.99 0.05 0.2831245 0.1550253 0.12424886 0.06007302
0.01 0.15 0.2961399 0.1186617 0.09774779 0.04138286
0.10 0.15 0.2905012 0.1278174 0.10876692 0.03876993
0.75 0.99 0.2817502 0.1770336 0.13952952 0.07232276
0.99 0.99 0.2817300 0.1796089 0.14184791 0.07272979
plot(z)
Figure 6: Fairness-accuracy tradeoff plot for COMPAS data, with misclassification rate (Y-axis) and correlation between predicted recidivism and race (X-axis).
The output of dsldFairUtils shows the misclassification rate across different unfairness levels, along with correlation values for each race. We observe a Fairness-Utility tradeoff: as the correlation between $S$ and predicted $Y$ increases, the overall misclassification rate decreases. Adjusting the unfairness parameter significantly impacts correlations, especially for African Americans, who show the widest range of changes. The plot visualizes the change in misclassification rate with correlation, treating unfairness as a 3rd hidden parameter.
We now look at the Law School Admissions dataset using dsldFrrm, with the several values of the unfairness parameter ranging from 0.01 to 0.99.
ex1 <- dsldUtilsFairness(law.school.admissions, lsat’, race1’, dsldFrrm’, unfairness = c(0.01, 0.05, 0.1, 0.25, 0.5, 0.75, 0.8, 0.9, 0.99), yesYVal = NULL, reps = 10)
print(ex1)
d MAPE asian black hisp other white
0.01 3.63402 0.0103227 0.0557155 0.0394955 0.00979468 0.0668236
0.05 3.55174 0.0399833 0.1344877 0.0680349 0.02208106 0.1533411
0.10 3.51030 0.0407389 0.1762693 0.0985790 0.02496056 0.1987062
0.25 3.42200 0.0602939 0.2680973 0.1509890 0.05899232 0.3095009
0.50 3.38134 0.0446075 0.3079764 0.1688408 0.06858589 0.3401874
0.99 3.38704 0.0344068 0.3049084 0.1876693 0.06376829 0.3422318
Figure 7: Fairness-accuracy tradeoff plot for LSA data, with MAPE (Y-axis) and correlation between predicted LSAT score and race (X-axis).
Similar to the COMPAS example, we observe a Fairness-Utility tradeoff. At high unfairness levels, the model achieves greater accuracy (MAPE = 3.38) but shows strong correlations (0.34) between predicted LSAT scores and some racial groups, particularly for White and African Americans. Prioritizing fairness with lower unfairness values slightly reduces accuracy but considerably reduces the respective correlations, especially for these subgroups. Figure 7 illustrates the tradeoff between MAPE and correlation across each race.
With dsld providing a range of fair machine learning functions, students can explore different modeling techniques and gain hands-on experience in developing responsible ML systems for real-world applications.
@@ 5 Discussion
In this paper, we introduce “Data Science Looks at Discrimination” (dsld) as a powerful tool for statistical education and machine learning through the lens of discrimination analysis and fair machine learning. The software includes analytical and graphical tools for detailed exploratory data analysis, allowing students to explore and visualize potential sources of bias. Additionally, the fair machine learning wrappers provide several fairness-constrained methods for easy deployment of fair ML algorithms. The accompanying Quarto book offers a comprehensive introduction to key statistical principles, using real-world examples to demonstrate dsld methods. It is accessible to students with just high school mathematics, equipping them with the knowledge and tools to conduct data analysis and apply machine learning systems responsibly.
Instructors and students are both encouraged to apply these methods to their own datasets for further analysis. The methods provided by the dsld package serve as a valuable addition to introductory statistics courses, helping students connect theoretical statistical concepts to practical use cases.
@@
The package can be accessed here: https://github.com/matloff/dsld/tree/master. All relevant datasets, functions, and implementation details can be found through the Github.